OpenStudy (anonymous):

WHAT VALUE MAKES C THE POLYNOMIAL A PERFECT SQUARE TRINOMIAL 4X2-32X+C

OpenStudy (mandre):

(2x-8)(2x-8) = 4x2 - 16x - 16x + 64 = 4x2 - 32x + 64 C = 64